A projectile is fired at an angle of 30 degrees above the horizontal from the top of a cliff 600 ft high. The initial speed of the projectile is 2000 ft/s

a. Find the range of projectile
b. Find the time of flight
1
Expert's answer
2020-09-24T11:09:22-0400

Let's first write the projections of initial velocity of the projectile on axis x- and y:


"v_{x} = v_0cos\theta,""v_y = v_0sin\theta,"

here, "v_0 = 2000 \ \dfrac{ft}{s}" is the initial velocity of the projectile, "\theta = 30^{\circ}" is the launch angle.

Let's write the equations of the projectile motion along axis "x"- and "y" (also, we choose the upwards as the positive direction):

"x=v_0tcos\theta,""y=v_0tsin\theta+ \dfrac{1}{2}gt^2,"

here, "x" is the range of the projectile, "y = -600 \ ft" is the vertical displacement of the projectile (or the height), "t" is the time of flight, "g = -32 \ \dfrac{ft}{s^2}" is the acceleration due to gravity.

We can find the time of flight from the second equation:


"-600 = 2000 \cdot sin30^{\circ} \cdot t - \dfrac{1}{2} \cdot 32 \cdot t^2,""16t^2 - 1000t -600 = 0."

This quadratic equation has two roots:


"t_1 = \dfrac{-(-1000)+ \sqrt{(-1000)^2-4 \cdot 16 \cdot (-600)}}{2 \cdot 16} = 63 \ s,""t_2 = \dfrac{-(-1000)- \sqrt{(-1000)^2-4 \cdot 16 \cdot (-600)}}{2 \cdot 16} = -0.6 \ s."

Since time can't be negative, the correct answer is "t = 63 \ s".

Finally, we can find the range of projectile from the first equation:


"x = 2000 \ \dfrac{ft}{s} \cdot cos30^{\circ} \cdot 63 \ s = 109116 \ ft."

Answer:

a) "x = 109116 \ ft."

b) "t = 63 \ s."